# Entropic uncertainty and coherence in Einstein-Gauss-Bonnet gravity

### v1

@@ I Introduction
The concept of higher-dimensional spacetime, motivated by heterotic string theory and braneworld models L3 ; L4 ; L5 ; L6 ; LL7 , extends general relativity beyond its conventional (3+1)-dimensional framework by introducing additional compact or extended dimensions. In the braneworld scenario, matter and gauge fields are confined to a (3+1)-dimensional brane, while gravity propagates in the higher-dimensional bulk, necessitating a generalized theory of gravity. Among such extensions, Lovelock gravity provides a natural and ghost-free framework L10 ; L11 ; L14 ; LL3 , with its second-order term yielding EGB gravity L15 ; L16 , which includes the cosmological constant, Einstein-Hilbert, and Gauss-Bonnet terms. The Gauss-Bonnet correction offers a nontrivial extension of Einstein gravity, leading to rich implications for black hole physics, thermodynamics, and holography. As four-dimensional general relativity remains incompatible with a complete quantum theory, exploring the quantum properties of higher-dimensional black holes in EGB gravity provides a promising route toward understanding the interplay between gravity and quantum effects L18 ; L19 ; L20 .
@@
Relativistic quantum information is an interdisciplinary field at the intersection of quantum information theory, quantum field theory, and general relativity R1 ; R2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 ; R32 ; R33 ; R34 ; R35 ; R36 ; R37 ; R38 ; R39 ; R40 ; R41 ; R42 ; R43 ; R44 ; R45 ; R46 ; R47 ; R48 ; R49 ; R50 ; RR51 ; R51 ; R52 ; ENM1 ; ENM2 . It has seen rapid theoretical and experimental development, offering novel insights into quantum phenomena in curved and Rindler spacetimes. In particular, quantum resources have been extensively explored in four-dimensional spacetime such as Schwarzschild black holes, Garfinkle-Horowitz-Strominger (GHS) dilaton black holes, and uniformly accelerated (noninertial) frames R1 ; R2 ; R3 ; R4 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 ; R32 ; R33 ; R34 . Despite substantial progress in related areas, the fundamental characteristics and distinctions between the entropic uncertainty relation (EUR) and quantum coherence in the context of EGB gravity remain unclear, particularly regarding their behavior for different quantum states. Since different quantum states may require different types of quantum resources, a thorough understanding of these features is essential for efficiently performing relativistic quantum information processing tasks. This gap in understanding serves as the key motivation for the present study. The uncertainty principle imposes fundamental limits on the simultaneous predictability of two non-commuting observables. To more accurately quantify this limitation, entropic formulations of the uncertainty principle were developed and subsequently extended to quantum-memory-assisted EUR, which reveal that quantum correlations can effectively reduce uncertainty R53 ; R54 . In parallel, quantum coherence viewed as a resource arising from basis-dependent superposition has also been found to contribute directly to entropic uncertainty R55 . Recently, the connection between quantum coherence and EUR has attracted attention and achieved some progress, but its underlying mechanism remains unclear R54 ; R56 . In this work, we focus on GHZ and W states as probes to systematically analyze the interplay between EUR and quantum coherence in the background of EGB gravity. By contrasting their distinct quantum informational responses to the gravitational environment, we aim to reveal how spacetime geometry influences fundamental quantum properties, thereby offering new insights into the physical implications of EGB gravity.
@@ III EUR and quantum coherence
This relation has been generalized in various directions. In particular, Boileau and Renes introduced the quantum-memory-assisted entropic uncertainty relation (QMA-EUR) for bipartite and tripartite systems M30 ; M31 . For the tripartite case, the relation takes the form
@@
respectively, where $\bm{\mathrm{I}}_{B(C)}$ is the identity operator. $S(X|B)$ is used to measure Bob’s uncertainty regarding Alice’s measurement result for the observable $X$. In simple terms, the uncertain relationship can be described by an uncertain game. Suppose there are three friends, namely Alice, Bob, and Charlie. Initially, they share a quantum state $\rho_{ABC}$. Next, Alice selects one of the two measurements to execute the command, either $X$ or $Z$, and informs her choice to Bob and Charlie who respectively hold the quantum memory $B$ and $C$. If she chooses $X$, Bob’s task is to predict Alice’s measurement result; if she chooses $Z$, then it is Charlie’s task to predict Alice’s measurement result. In addition, the QMA-EUR has attracted considerable attention due to its wide range of applications, including quantum metrology and entanglement detection R53 . Experimentally, significant progress has been made in verifying and applying QMA-EUR across different platforms M46 ; M48 ; M51 .
@@ IV.1 GHZ state
$U_{AB_{\mathrm{I}}C_{\mathrm{I}}}^{GHZ}=\frac{-\kappa\ln\kappa+\kappa\ln[(2(\kappa+1)]+\ln(2\kappa+2)+\delta\ln 2}{\ln 4},$ (32)
@@
$U_{A_{\mathrm{I}}BC}^{GHZ}=\frac{-\cos^{2}\zeta\ln\frac{\cos^{2}\zeta}{2}-\sin^{2}\zeta\ln\frac{\sin^{2}\zeta}{2}+\ln 2}{\ln 4}.$ (37)
@@
In Fig.3, the measurement uncertainty and quantum coherence of the GHZ state in the fermionic field are shown as functions of the Gauss-Bonnet parameter $\alpha$ and the horizon radius $r_{h}$ for two physical scenarios in spacetime dimensions $d=5,6,7$ with fixed frequency $\omega=1$. From Fig.3(a), we observe that the measurement uncertainty in both scenarios decreases monotonically with increasing $\alpha$. This behavior suggests that in higher-dimensional EGB spacetime, a larger Gauss-Bonnet coupling enhances the rigidity of spacetime, suppresses quantum fluctuations, and consequently reduces measurement uncertainty. Consistently, Fig.3(c) and (e) show that both the $l_{1}$-norm of coherence and REC increase monotonically with $\alpha$. Fig.3(b) shows that for $d=6,7$, the uncertainty decreases monotonically with increasing horizon radius $r_{h}$. This can be attributed to the fact that a larger $r_{h}$ corresponds to a lower Hawking temperature $T$, leading to weaker quantum fluctuations and a more rigid spacetime. In contrast, for $d=5$, the uncertainty exhibits a non-monotonic behavior first increasing and then decreasing arising from the peculiar thermodynamic properties of the five-dimensional EGB spacetime as determined by Eqs.(10) and (11). Consistently, Fig.3(d) and (f) illustrate that for $d=6,7$, quantum coherence grows with $r_{h}$, since the enlarged horizon suppresses decoherence induced by spacetime curvature and thermal noise. For $d=5$, however, quantum coherence displays a non-monotonic dependence on $r_{h}$, first decreasing and then increasing, which reflects the distinctive role of the five-dimensional spacetime structure.
@@
From Fig.3, we find that the overall trends of measurement uncertainty and coherence with respect to the black hole parameters are completely opposite. This contrast arises from the fact that quantum coherence embodies a form of nonclassical correlation: stronger coherence corresponds to reduced measurement uncertainty, and conversely, weaker coherence leads to higher measurement uncertainty. Moreover, we also find that the gravitational effects in Scenario 1 exert a stronger influence on both measurement uncertainty and coherence of GHZ state compared to those in Scenario 2.

### v2

@@ I Introduction
The concept of higher-dimensional spacetime, motivated by heterotic string theory and braneworld models L3 ; L4 ; L5 ; L6 ; LL7 , extends general relativity beyond its conventional (3+1)-dimensional framework by introducing additional compact or extended dimensions. In the braneworld scenario, matter and gauge fields are confined to a (3+1)-dimensional brane, while gravity propagates in the higher-dimensional bulk, necessitating a generalized theory of gravity. Among such extensions, Lovelock gravity provides a natural and ghost-free framework L10 ; L11 ; L14 ; LL3 , with its second-order term yielding EGB gravity L15 ; L16 , which includes the cosmological constant, Einstein-Hilbert, and Gauss-Bonnet terms. The Gauss-Bonnet correction offers a nontrivial extension of Einstein gravity, leading to rich implications for black hole physics, thermodynamics, and holography. As four-dimensional general relativity remains incompatible with a complete quantum theory, exploring the quantum properties of higher-dimensional black holes in EGB gravity provides a promising route toward understanding the interplay between gravity and quantum effects L18 ; L19 ; L20 ; LL20 .
@@
Relativistic quantum information is an interdisciplinary field at the intersection of quantum information theory, quantum field theory, and general relativity R1 ; RR1 ; R2 ; R3 ; R4 ; RR2 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 ; R32 ; R33 ; R34 ; R35 ; R36 ; R37 ; R38 ; R39 ; R40 ; R41 ; R42 ; R43 ; ENM5 ; R44 ; R45 ; R46 ; R47 ; R48 ; R49 ; R50 ; RR51 ; R51 ; R52 ; ENM1 ; ENM2 ; ENM4 ; ENM6 ; ENM7 ; ENM8 ; ENM9 ; AVM1 ; AVM2 . It has seen rapid theoretical and experimental development, offering novel insights into quantum phenomena in curved and Rindler spacetimes. In particular, quantum resources have been extensively explored in four-dimensional spacetime such as Schwarzschild black holes, Garfinkle-Horowitz-Strominger (GHS) dilaton black holes, and uniformly accelerated (noninertial) frames R1 ; RR1 ; R2 ; R3 ; R4 ; RR2 ; R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R18 ; R19 ; R20 ; R21 ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R29 ; R30 ; R31 ; R32 ; R33 ; R34 . Despite substantial progress in related areas, the fundamental characteristics and distinctions between the entropic uncertainty relation (EUR) and quantum coherence in the context of EGB gravity remain unclear, particularly regarding their behavior for different quantum states. Since different quantum states may require different types of quantum resources, a thorough understanding of these features is essential for efficiently performing relativistic quantum information processing tasks. This gap in understanding serves as the key motivation for the present study. The uncertainty principle imposes fundamental limits on the simultaneous predictability of two non-commuting observables. To more accurately quantify this limitation, entropic formulations of the uncertainty principle were developed and subsequently extended to quantum-memory-assisted EUR, which reveal that quantum correlations can effectively reduce uncertainty R53 ; R54 . In parallel, quantum coherence viewed as a resource arising from basis-dependent superposition has also been found to contribute directly to entropic uncertainty R55 . Recently, the connection between quantum coherence and EUR has attracted attention and achieved some progress, but its underlying mechanism remains unclear R54 ; R56 . In this work, we focus on GHZ and W states as probes to systematically analyze the interplay between EUR and quantum coherence in the background of EGB gravity. By contrasting their distinct quantum informational responses to the gravitational environment, we aim to reveal how spacetime geometry influences fundamental quantum properties, thereby offering new insights into the physical implications of EGB gravity.
@@ III EUR and quantum coherence
This relation has been generalized in various directions. In particular, Boileau and Renes introduced the quantum-memory-assisted entropic uncertainty relation (QMA-EUR) for bipartite and tripartite systems M32 ; M33 ; M30 ; M31 ; M34 ; M35 . For the tripartite case, the relation takes the form
@@
respectively, where $\bm{\mathrm{I}}_{B(C)}$ is the identity operator. $S(X|B)$ is used to measure Bob’s uncertainty regarding Alice’s measurement result for the observable $X$. In simple terms, the uncertain relationship can be described by an uncertain game. Suppose there are three friends, namely Alice, Bob, and Charlie. Initially, they share a quantum state $\rho_{ABC}$. Next, Alice selects one of the two measurements to execute the command, either $X$ or $Z$, and informs her choice to Bob and Charlie who respectively hold the quantum memory $B$ and $C$. If she chooses $X$, Bob’s task is to predict Alice’s measurement result; if she chooses $Z$, then it is Charlie’s task to predict Alice’s measurement result. In addition, the QMA-EUR has attracted considerable attention due to its wide range of applications, including quantum metrology and entanglement detection R53 . Experimentally, significant progress has been made in verifying and applying QMA-EUR across different platforms M46 ; MM51 ; M48 ; MM52 ; M51 .
@@ IV.1 GHZ state
$U_{AB_{\mathrm{I}}C_{\mathrm{I}}}^{GHZ}=\frac{1}{2}[-\kappa\log_{2}\kappa+\kappa\log_{2}[(2(\kappa+1)]+\log_{2}(2\kappa+2)+\delta],$ (32)
@@
$U_{A_{\mathrm{I}}BC}^{GHZ}=1+\frac{1}{2}[-\cos^{2}\zeta\log_{2}(\cos^{2}\zeta)-\sin^{2}\zeta\log_{2}(\sin^{2}\zeta)].$ (37)
@@
In Fig.3, the measurement uncertainty and quantum coherence of the GHZ state in the fermionic field are shown as functions of the Gauss-Bonnet parameter $\alpha$ and the horizon radius $r_{h}$ for two physical scenarios in spacetime dimensions $d=5,6,7$ with fixed frequency $\omega=1$. From Fig.3(a), we observe that the measurement uncertainty in both scenarios decreases monotonically with increasing $\alpha$. This behavior suggests that in higher-dimensional EGB spacetime, a larger Gauss-Bonnet coupling enhances the rigidity of spacetime, suppresses quantum fluctuations, and consequently reduces measurement uncertainty. Consistently, Fig.3(c) and (e) show that both the $l_{1}$-norm of coherence and REC increase monotonically with $\alpha$. Fig.3(b) shows that for $d=6,7$, the uncertainty decreases monotonically with increasing horizon radius $r_{h}$. This can be attributed to the fact that a larger $r_{h}$ corresponds to a lower Hawking temperature $T$, leading to weaker quantum fluctuations and a more rigid spacetime. In contrast, for $d=5$, the uncertainty exhibits a non-monotonic behavior first increasing and then decreasing arising from the peculiar thermodynamic properties of the five-dimensional EGB spacetime as determined by Eqs.(10) and (11). Consistently, Fig.3(d) and (f) illustrate that for $d=6,7$, quantum coherence grows with $r_{h}$, since the enlarged horizon suppresses decoherence induced by spacetime curvature and thermal noise. For $d=5$, however, quantum coherence displays a non-monotonic dependence on $r_{h}$, first decreasing and then increasing, which reflects the distinctive role of the five-dimensional spacetime structure. The differences in the behavior of QMA-EUR and quantum coherence between $d=5$ and $d>5$ dimensions are rooted in the distinct thermodynamic properties of EGB gravity in higher dimensions. Specifically, in $d>5$, the measurement uncertainty decreases monotonically with increasing event horizon radius, while quantum coherence increases, reflecting a more stable thermodynamic state. However, in $d=5$, both the measurement uncertainty and coherence exhibit non-monotonic behavior, which can be attributed to the unique thermodynamic properties of the five-dimensional EGB black hole. Incorporating these observations into the analysis of entropic uncertainty and quantum coherence, we show that the presence of distinctive thermodynamic properties in $d=5$ dimensions leads to non-monotonic changes in these quantities. This highlights the importance of the specific dimensional characteristics of the spacetime when considering quantum informational resources in the context of EGB gravity. This explanation clarifies the reason for the dimensional dependence and provides insight into the interplay between spacetime geometry, quantum fluctuations, and gravitational decoherence, which are crucial for understanding the behavior of quantum resources in curved spacetime.
@@
From Fig.3, we find that the overall trends of measurement uncertainty and coherence with respect to the black hole parameters are completely opposite. This contrast arises from the fact that quantum coherence embodies a form of nonclassical correlation: stronger coherence corresponds to reduced measurement uncertainty, and conversely, weaker coherence leads to higher measurement uncertainty. The opposite variation trends of QMA-EUR and quantum coherence in the EGB spacetime, which both depend on the Gauss-Bonnet coupling and the event horizon radius, can be explained through their relationship with the thermodynamic properties of the black hole in higher dimensions. In higher dimensions, the Gauss-Bonnet coupling $\alpha$ modifies spacetime curvature, influencing thermodynamic quantities such as temperature. An increase in Gauss-Bonnet coupling $\alpha$ or the event horizon radius $r_{h}$ generally enhances thermodynamic stability. This stability suppresses quantum fluctuations, thereby reducing entropic uncertainty. Conversely, greater stability helps preserve quantum correlations, leading to enhanced coherence. Moreover, we also find that the gravitational effects in Scenario 1 exert a stronger influence on both measurement uncertainty and coherence of GHZ state compared to those in Scenario 2.
@@ V Conclusions
Although the present work is primarily theoretical, recent advances in experimental quantum technologies suggest that the physical scenarios considered here may be explored in realistic or analog settings. Quantum optical experiments using the Micius satellite have already tested quantum field behavior in curved spacetime by transmitting entangled photon pairs through different regions of Earth’s gravitational potential, offering a direct experimental platform to investigate gravity induced decoherence and non-unitary effects qhx7 . Likewise, atomic clock experiments have measured gravitational redshifts within millimeter-scale samples of ultracold strontium atoms, achieving sensitivities relevant for probing quantum coherence at distinct gravitational potentials qhx8 . In parallel, the development of compact quantum microsatellites and satellite-based quantum key distribution networks shows that quantum coherence and entanglement can be distributed across large gravitational gradients between ground and orbit qhx9 . Taken together, these developments indicate that analog implementations of relativistic quantum information processes-such as qubits subjected to varying gravitational fields or near an event horizon analog potentials-are increasingly becoming experimentally accessible. The theoretical framework presented in this work therefore provides a foundation for interpreting these emerging experiments and for guiding the design of future analog gravity platforms capable of simulating qubit dynamics and coherence redistribution in curved spacetime.